# Risk-Averse Constrained Reinforcement Learning with Optimized Certainty Equivalents

### v1

@@ 1 Introduction
Autonomous agents are often used in settings where they must handle several conflicting requirements while maximizing a main objective, such as navigating a maze without hitting any walls. These conflicting requirements are often modeled with constrained reinforcement learning (RL) (see Sutton and Barto [42]) and are either solved as a multi-objective problem with conflicting requirements controlled by weights, chosen manually or through hyperparameter tuning (e.g., see Borkar [10], Achiam et al. [2], Mannor and Shimkin [28], Tessler et al. [45], Mania et al. [27]), or by using primal-dual methods which optimize over the weights [45, 5]. Others have approached constrained RL through reward-free techniques, e.g., [30], or strategic exploration [54]. The work of Paternain et al. [33] has given theoretical support to primal-dual approaches by proving that, despite their inherent nonconvexity, constrained RL models exhibit zero duality gap under mild assumptions. However, in the standard (risk-neutral) setting (including that of Paternain et al. [33]), the objectives and constraints are often formulated as expected values of a (discounted) sum of rewards. Nonetheless, in many practical scenarios, especially in the context of safety [24] and other high-stakes applications, the standard expectation may not be strict enough to describe the desired behavior. For example, instead of maximizing the on-average return (e.g., of an investment), one may want to learn policies that mitigate risk (e.g., of losing money).
@@
Risk measures are often used to quantify this risk in problems with uncertain (random) outcomes (see, e.g., Shapiro et al. [41]). More broadly, stochastic optimization with risk measures in the objectives and/or constraints has been studied in the online setting [26], bandits [11, 37, 44], statistical learning [3, 15, 32, 48], non-convex resource allocation problems [21], and others [20, 55]. These methods and analyses are not readily applicable in our setting either due to non-convexity/concavity in the objective and constraints or due to the fact that the risk envelope for our problem has dependence on the policy $\pi$.
@@
In reinforcement learning specifically, there is a deep literature on risk-averse methods [31, 53, 13, 14, 29, 35, 49, 51, 50] with many applications in the area of robust control [52, 23]. While the most conservative approach consists of learning policies robust to worst-case scenarios [1, 18], we often prefer to take into account a set of events with significant probability. To do this, Huang et al. [17], Tamar et al. [43] consider the policy gradient method under coherent risk measures, where the risk measure replaces the expectation over the discounted sum of rewards. The work of Chow et al. [12] handles similar CVaR risk objectives with CVaR risk constraints by utilizing Lagrangian relaxation. Xu et al. [51], Wang et al. [50] study risk-sensitive RL with OCEs. Recently, Bonetti et al. [7] propose an alternative risk-aware framework for (unconstrained) reinforcement learning which places the risk measure in the occupancy measure to derive what the authors call the reward-based conditional value at risk (RCVaR) and mean-mean absolute deviation (Mean-RMAD), in contrast to the standard return-based formulations (like that of [12, 17, 43]).
@@
  • We extend the setting of Bonetti et al. [7] and, to the best of our knowledge, our work is the first to handle reward-based constraints, covering a large class of risk measures (OCEs);
@@
The setting of Bonetti et al. [7] is closely related to this work but is restricted to the unconstrained case, while our setting handles reward-based risk constraints as well. Our algorithm differs from the methodology of [7], which is based on a block-coordinate descent scheme, the analysis of which does not readily extend to the type of optimization problems arising in the constrained setting. Further, [7] requires exact policy solvers but we work with inexact ones (cf. Assumption 3.6). Compared to other risk-averse constrained RL methods, our partial Lagrangian relaxation is exact under a certain constraint qualification, whereas no meaningful (let alone exact) relation between the original primal problem and the employed Lagrangian relaxation has been shown in other works (e.g., [12]).
@@ 2.1 Risk Measures
In the context of stochastic optimization, is it well-known that expectations are unable to capture “risky" events (related to statistical variability, dispersion, or fat-tail behavior of the randomness). Risk-averse optimization aims to minimize the risk associated with such events, which is captured by certain functionals known as risk measures. For an extended discussion on risk-averse optimization, we refer the reader to Shapiro et al. [41, Chapter 6].
@@
If $Z$ represents rewards (as often in RL), we consider risk measures in the supremal convolution form, i.e., $\widetilde{\rho}(Z)=\sup_{t}\{t+\mathbb{E}\left[g(Z-t)\right]\}$. For example, we can consider the reflected $\mathrm{CVaR}^{\beta}(-Z)$ (for reward maximization) by setting $g(u)=-\frac{1}{\beta}(-u)_{+}$. In this notation, $g$ is called a utility function. Under certain conditions on the utility (e.g., satisfied by $g(u)=-(-u)_{+}$), these risk measures are known as optimized certainty equivalents (OCEs) (see Ben-Tal and Teboulle [4, Definition 2.2], where the reader can find many practical examples and properties). The class of OCEs is very rich (including many risk measures of practical value) and satisfies several important properties.
@@ 2.2 Risk-Neutral RL
where $(s_{\tau}^{\pi},a_{\tau}^{\pi})$ denotes the state-action vector evolving under policy $\pi$. It is well-known that the above can be reformulated in terms of the (discounted) occupancy (or occupation) measure of a policy, defined as the discounted mixture $\mathrm{d}\nu^{\pi}(\cdot,\bullet)=(1-\gamma)\sum_{\tau=0}^{\infty}\gamma^{\tau}\mathrm{d}p_{\pi}^{\tau}(\cdot,\bullet)$, where $p_{\pi}^{\tau}$ denotes the Borel probability measure induced by vector $(s_{\tau}^{\pi},a_{\tau}^{\pi})$. Let $\mathcal{R}$ denote the convex and weakly compact set of all occupation measures [8] induced by the policies $\pi\in\mathcal{P(S)}$. We can then rewrite the RL problem as follows.
Problem (Convex Analytic Dual Form [^2] This notion of casting dynamic optimization problems into abstract “static” optimization problems over a closed convex set of measures (as above) is referred to as the convex analytic approach by Borkar [9]. Thus, we will refer to these equivalent problems (1) and (2) as the “primal” and “convex analytic dual” formulations (not to be confused with duality in the Lagrangian sense) throughout.).
@@
It is known that the occupation measure $\nu^{\pi}$ is in a one-to-one relationship with a policy $\pi$, i.e., if two policies have the same occupation measure, they must be the same (Sutton and Barto [42]).
@@ 2.3 Risk-Averse RL Formulations
There are several different frameworks for modeling “risk-awareness” and “safety” in RL. Existing works which explore risk-averse RL formulations have often taken the approach of replacing the expectations in the “primal” formulation (1) with risk measures, such as in the studies of Huang et al. [17], Tamar et al. [43], Chow et al. [12], among others:
@@
Alternatively, one could generalize the tower property of expectations and iteratively evaluate a risk measure at each decision stage (e.g., see Ruszczyński [36]). A less common formulation –of interest herein– applies a risk measure on the occupancy measure, rather than the original probability space, as in Bonetti et al. [7]:
@@
This relationship is summarized in Table 1. As one can see, the risk-averse problem involves a maximization over an additional variable $t$ which depends on $\pi$ (in this sequence). Importantly, given a fixed $t$, maximization over $\pi$ resembles a problem of the form of (1) with a modified reward function. Thus one can take advantage of several existing algorithms which solve (1). This has been explored in the unconstrained setting of Bonetti et al. [7, Algorithm 2].
@@ 2.4 Constraints
Here, we highlight that Bonetti et al. [7] do not consider the constrained RL setting (which requires a dedicated analysis of the relation between the Lagrangian relaxation and the original constrained problem), while others using return-based risk objectives and constraints in the convex analytic “primal” space (e.g., Chow et al. [12]) are not able to show any meaningful relation between the primal problem and their employed Lagrangian relaxation.
@@ 3.1 Problem Description and Equivalent Reformulations
Consider any OCE, and assume that $r_{i}$ are bounded for all $i\in\{0\}\cup[m].$ Then, by Ben-Tal and Teboulle [4, Proposition 2.1], we have that
@@
The proof is based on Chow et al. [12], and given in Appendix A.2. Although it may seem somewhat redundant, (7) is a maximization problem jointly over the variables $\pi,t$, subject to functional inequality constraints. Thus, the Lagrangian associated to (7) reads ${\mathcal{L}}(\pi,t,\lambda)\coloneqq{\mathbb{E}}[\widehat{{\mathcal{L}}}(\pi,t,\lambda)]$, with
@@
Then, the primal problem is $P^{*}=\sup_{\pi,t}\inf_{\lambda\geq 0}{\mathcal{L}}(\pi,t,\lambda)$. Note that the above, for fixed $t,\lambda$, is the Lagrangian for a risk-neutral constrained problem (similar to those considered by Paternain et al. [33]), parameterized by $t$, which controls risk-aversion (in harmony with the choice of $\beta$).
@@ 3.2 Partial Lagrangian Relaxation
In order to solve the (inner) policy optimization problem, we parametrize the policy by a vector $\theta\in\Theta\subset\mathbb{R}^{p}$, representing the coefficients of, say, a neural network, assuming that $\Theta$ is a compact set (which is the case in practice). We have shown that under Assumption 3.4, the proposed partial Lagrangian relaxation is exact. We argue that this remains almost true (in the sense described below) even if we utilize parametrized policies. In what follows, we assume that $\pi_{\theta}$ is an $\epsilon$-universal parametrization of measures in $\mathcal{P}(\mathcal{S})$, according to Paternain et al. [33, Definition 1].
@@ 3.3 Algorithm and Analysis
We prove Lemma 3.7 (which readily applies to parametrized policies as well) in Appendix A.6. Following Lin et al. [25], we analyze our method under the assumption of Lipschitz smoothness of the function $\mathcal{L}(\pi_{\theta^{*}(t,\lambda)},t,\lambda)$, for $\theta^{*}(t,\lambda)\in\arg\max_{\theta\in\Theta}\mathcal{L}(\pi_{\theta},t,\lambda)$, which we state below.
@@
As appears in prior work [25], we make the standard (and, in our case, mild) assumption that the exact stochastic (sub)gradient oracles are unbiased and have at most $\sigma^{2}$ variance.
@@
Note that this result says we only need single trajectories ($n=1$). Thus our algorithm can be run online. The proof, which is detailed in Appendix A.8, extends the developments of [25], by incorporating biased (sub)gradient samples, occurring due to the use of inexact oracles. If $\delta=\mathcal{O}(\epsilon^{2})$, then we recover an $\epsilon$-stationary point and the complexity of [25].
@@ 4 Experiments
We conduct experiments on locomotion tasks from the Safety-Gymnasium benchmark [19].
@@ 4.1 Setup
We consider two constraints in the environment: safe navigation and safe velocity. The first involves simple problems with low-dimensional state and action spaces, where the agent must navigate under rules (e.g., along a path or without hitting obstacles). The second is more challenging, involving standard high-dimensional control tasks based on MuJoCo-v4 agents [46]. At each timestep, the agent receives a cost of +1 if it violates a constraint and 0 otherwise. These violations do not affect environment dynamics—they neither terminate episodes nor alter transitions. Importantly, our experiments do not access this cost directly during training; instead, the agent manages risk by being risk-aware in its velocities.
@@
Proximal Policy Optimization (PPO) [39] is used as a black-box solver to optimize the policy $\pi$, implementing line 4 of Algorithm 1. The dual variable $\lambda$ and auxiliary $t$ are updated via stochastic gradient steps following the rest of Algorithm 1. For further experimental details, see Appendix B.
@@
(a) Environment: HalfCheetah   $c=1.450$   $\beta$-upper quantile: $1.419$   Converged $t$-value: $1.417$
(b) Environment: Hopper   $c=0.373$   $\beta$-upper quantile: $0.370$   Converged $t$-value: $0.370$
(c) Environment: Swimmer   $c=0.228$   $\beta$-upper quantile: $0.248$   Converged $t$-value: $0.207$
(d) Environment: Walker2d   $c=1.171$   $\beta$-upper quantile: $1.133$   Converged $t$-value: $1.122$

### v2

@@ 1 Introduction
Autonomous agents are often used in settings where they must handle several conflicting requirements while maximizing a main objective, such as navigating a maze without hitting any walls. These conflicting requirements are often modeled with constrained reinforcement learning (RL) (see sutton_rl_textbook) and are either solved as a multi-objective problem with conflicting requirements controlled by weights, chosen manually or through hyperparameter tuning (e.g., see borkar_actor_critic, constr_policy_opt, multi_criterion_rl, reward_constr_policy_opt, random_search_static_linear_rl), or by using primal-dual methods which optimize over the weights [reward_constr_policy_opt, online_actor_critic_alg_constr_rl]. Others have approached constrained RL through reward-free techniques, e.g., [pmlr-v162-miryoosefi22a], or strategic exploration [pmlr-v139-yu21b]. The work of constrained_rl_zero_duality_gap has given theoretical support to primal-dual approaches by proving that, despite their inherent nonconvexity, constrained RL models exhibit zero duality gap under mild assumptions. However, in the standard (risk-neutral) setting (including that of constrained_rl_zero_duality_gap), the objectives and constraints are often formulated as expected values of a (discounted) sum of rewards. Nonetheless, in many practical scenarios, especially in the context of safety [ai_safety_gridworlds] and other high-stakes applications, the standard expectation may not be strict enough to describe the desired behavior. For example, instead of maximizing the on-average return (e.g., of an investment), one may want to learn policies that mitigate risk (e.g., of losing money).
@@
Risk measures are often used to quantify this risk in problems with uncertain (random) outcomes (see, e.g., stoch_programming_text_shapiro). More broadly, stochastic optimization with risk measures in the objectives and/or constraints has been studied in the online setting [stoch_primal_dual_cvar], bandits [pmlr-v89-cardoso19a, NIPS2012_83f25503, seq_dec_coherent_risk], statistical learning [bedi2020nonparametric, svm_convex_risk_func, JMLR:v18:15-566, Vitt2018RiskAverseC], non-convex resource allocation problems [risk_constrained_resource_alloc], and others [Jiang2018554, approx_risk_submodular]. These methods and analyses are not readily applicable in our setting either due to non-convexity/concavity in the objective and constraints or due to the fact that the risk envelope for our problem has dependence on the policy $\pi$.
@@
In reinforcement learning specifically, there is a deep literature on risk-averse methods [parallel_nonstat_direct_policy, 8247278, NEURIPS2020_fdc42b6b, geibel2005risk, mihatsch2002risk, PrashanthL2018RiskSensitiveRL, 2023_wang_near_minimax_cvar, 2023_xu_regret_recursive_oce, wang2025reductionsapproachrisksensitivereinforcement] with many applications in the area of robust control [10.3389/frobt.2021.617839, lam2023riskaware]. While the most conservative approach consists of learning policies robust to worst-case scenarios [abdullah2019wasserstein, risk_aware_q_learning_mdp], we often prefer to take into account a set of events with significant probability. To do this, huang2021convergence, NIPS2015_024d7f84 consider the policy gradient method under coherent risk measures, where the risk measure replaces the expectation over the discounted sum of rewards. The work of risk_contr_rl_percentile_risk handles similar CVaR risk objectives with CVaR risk constraints by utilizing Lagrangian relaxation. 2023_xu_regret_recursive_oce, wang2025reductionsapproachrisksensitivereinforcement study risk-sensitive RL with OCEs. Recently, bonetti_risk_averse_rl_coherent_risk propose an alternative risk-aware framework for (unconstrained) reinforcement learning which places the risk measure in the occupancy measure to derive what the authors call the reward-based conditional value at risk (RCVaR) and mean-mean absolute deviation (Mean-RMAD), in contrast to the standard return-based formulations (like that of [risk_contr_rl_percentile_risk, huang2021convergence, NIPS2015_024d7f84]).
@@
  • We extend the setting of bonetti_risk_averse_rl_coherent_risk and, to the best of our knowledge, our work is the first to handle reward-based constraints, covering a large class of risk measures (OCEs);
@@
The setting of bonetti_risk_averse_rl_coherent_risk is closely related to this work but is restricted to the unconstrained case, while our setting handles reward-based risk constraints as well. Our algorithm differs from the methodology of [bonetti_risk_averse_rl_coherent_risk], which is based on a block-coordinate descent scheme, the analysis of which does not readily extend to the type of optimization problems arising in the constrained setting. Further, [bonetti_risk_averse_rl_coherent_risk] requires exact policy solvers but we work with inexact ones (cf. Assumption 3.6). Compared to other risk-averse constrained RL methods, our partial Lagrangian relaxation is exact under a certain constraint qualification, whereas no meaningful (let alone exact) relation between the original primal problem and the employed Lagrangian relaxation has been shown in other works (e.g., [risk_contr_rl_percentile_risk]).
@@ 2.1 Risk Measures
In the context of stochastic optimization, is it well-known that expectations are unable to capture “risky" events (related to statistical variability, dispersion, or fat-tail behavior of the randomness). Risk-averse optimization aims to minimize the risk associated with such events, which is captured by certain functionals known as risk measures. For an extended discussion on risk-averse optimization, we refer the reader to stoch_programming_text_shapiro.
@@
If $Z$ represents rewards (as often in RL), we consider risk measures in the supremal convolution form, i.e., $\widetilde{\rho}(Z)=\sup_{t}\{t+\mathbb{E}\left[g(Z-t)\right]\}$. For example, we can consider the reflected $\mathrm{CVaR}^{\beta}(-Z)$ (for reward maximization) by setting $g(u)=-\frac{1}{\beta}(-u)_{+}$. In this notation, $g$ is called a utility function. Under certain conditions on the utility (e.g., satisfied by $g(u)=-(-u)_{+}$), these risk measures are known as optimized certainty equivalents (OCEs) (see ben-tal_teboulle_oce_07, where the reader can find many practical examples and properties). The class of OCEs is very rich (including many risk measures of practical value) and satisfies several important properties.
@@ 2.2 Risk-Neutral RL
where $(s_{\tau}^{\pi},a_{\tau}^{\pi})$ denotes the state-action vector evolving under policy $\pi$. It is well-known that the above can be reformulated in terms of the (discounted) occupancy (or occupation) measure of a policy, defined as the discounted mixture $\mathrm{d}\nu^{\pi}(\cdot,\bullet)=(1-\gamma)\sum_{\tau=0}^{\infty}\gamma^{\tau}\mathrm{d}p_{\pi}^{\tau}(\cdot,\bullet)$, where $p_{\pi}^{\tau}$ denotes the Borel probability measure induced by vector $(s_{\tau}^{\pi},a_{\tau}^{\pi})$. Let $\mathcal{R}$ denote the convex and weakly compact set of all occupation measures [Borkar_convex_analytic_journal] induced by the policies $\pi\in\mathcal{P(S)}$. We can then rewrite the RL problem as follows.
Problem (Convex Analytic Dual Form [^2] This notion of casting dynamic optimization problems into abstract “static” optimization problems over a closed convex set of measures (as above) is referred to as the convex analytic approach by Borkar2002_convex_analytic_mdp. Thus, we will refer to these equivalent problems (1) and (2) as the “primal” and “convex analytic dual” formulations (not to be confused with duality in the Lagrangian sense) throughout.).
@@
It is known that the occupation measure $\nu^{\pi}$ is in a one-to-one relationship with a policy $\pi$, i.e., if two policies have the same occupation measure, they must be the same (sutton_rl_textbook).
@@ 2.3 Risk-Averse RL Formulations
There are several different frameworks for modeling “risk-awareness” and “safety” in RL. Existing works which explore risk-averse RL formulations have often taken the approach of replacing the expectations in the “primal” formulation (1) with risk measures, such as in the studies of huang2021convergence, NIPS2015_024d7f84, risk_contr_rl_percentile_risk, among others:
@@
Alternatively, one could generalize the tower property of expectations and iteratively evaluate a risk measure at each decision stage (e.g., see ruszczynski2010risk). A less common formulation –of interest herein– applies a risk measure on the occupancy measure, rather than the original probability space, as in bonetti_risk_averse_rl_coherent_risk:
@@
This relationship is summarized in Table 1. As one can see, the risk-averse problem involves a maximization over an additional variable $t$ which depends on $\pi$ (in this sequence). Importantly, given a fixed $t$, maximization over $\pi$ resembles a problem of the form of (1) with a modified reward function. Thus one can take advantage of several existing algorithms which solve (1). This has been explored in the unconstrained setting of bonetti_risk_averse_rl_coherent_risk.
@@ 2.4 Constraints
Here, we highlight that bonetti_risk_averse_rl_coherent_risk do not consider the constrained RL setting (which requires a dedicated analysis of the relation between the Lagrangian relaxation and the original constrained problem), while others using return-based risk objectives and constraints in the convex analytic “primal” space (e.g., risk_contr_rl_percentile_risk) are not able to show any meaningful relation between the primal problem and their employed Lagrangian relaxation.
@@ 3.1 Problem Description and Equivalent Reformulations
Consider any OCE, and assume that $r_{i}$ are bounded for all $i\in\{0\}\cup[m].$ Then, by ben-tal_teboulle_oce_07, we have that
@@
The proof is based on risk_contr_rl_percentile_risk, and given in Appendix A.2. Although it may seem somewhat redundant, (7) is a maximization problem jointly over the variables $\pi,t$, subject to functional inequality constraints. Thus, the Lagrangian associated to (7) reads ${\mathcal{L}}(\pi,t,\lambda)\coloneqq{\mathbb{E}}[\widehat{{\mathcal{L}}}(\pi,t,\lambda)]$, with
@@
Then, the primal problem is $P^{*}=\sup_{\pi,t}\inf_{\lambda\geq 0}{\mathcal{L}}(\pi,t,\lambda)$. Note that the above, for fixed $t,\lambda$, is the Lagrangian for a risk-neutral constrained problem (similar to those considered by constrained_rl_zero_duality_gap), parameterized by $t$, which controls risk-aversion (in harmony with the choice of $\beta$).
@@ 3.2 Partial Lagrangian Relaxation
In order to solve the (inner) policy optimization problem, we parametrize the policy by a vector $\theta\in\Theta\subset\mathbb{R}^{p}$, representing the coefficients of, say, a neural network, assuming that $\Theta$ is a compact set (which is the case in practice). We have shown that under Assumption 3.4, the proposed partial Lagrangian relaxation is exact. We argue that this remains almost true (in the sense described below) even if we utilize parametrized policies. In what follows, we assume that $\pi_{\theta}$ is an $\epsilon$-universal parametrization of measures in $\mathcal{P}(\mathcal{S})$, according to constrained_rl_zero_duality_gap.
@@ 3.3 Algorithm and Analysis
We prove Lemma 3.7 (which readily applies to parametrized policies as well) in Appendix A.6. Following pmlr-v119-lin20a, we analyze our method under the assumption of Lipschitz smoothness of the function $\mathcal{L}(\pi_{\theta^{*}(t,\lambda)},t,\lambda)$, for $\theta^{*}(t,\lambda)\in\arg\max_{\theta\in\Theta}\mathcal{L}(\pi_{\theta},t,\lambda)$, which we state below.
@@
As appears in prior work [pmlr-v119-lin20a], we make the standard (and, in our case, mild) assumption that the exact stochastic (sub)gradient oracles are unbiased and have at most $\sigma^{2}$ variance.
@@
Note that this result says we only need single trajectories ($n=1$). Thus our algorithm can be run online. The proof, which is detailed in Appendix A.8, extends the developments of [pmlr-v119-lin20a], by incorporating biased (sub)gradient samples, occurring due to the use of inexact oracles. If $\delta=\mathcal{O}(\epsilon^{2})$, then we recover an $\epsilon$-stationary point and the complexity of [pmlr-v119-lin20a].
@@ 4 Experiments
We conduct experiments on locomotion tasks from the Safety-Gymnasium benchmark [safety_gym].
@@ 4.1 Setup
We consider two constraints in the environment: safe navigation and safe velocity. The first involves simple problems with low-dimensional state and action spaces, where the agent must navigate under rules (e.g., along a path or without hitting obstacles). The second is more challenging, involving standard high-dimensional control tasks based on MuJoCo-v4 agents [mujoco]. At each timestep, the agent receives a cost of +1 if it violates a constraint and 0 otherwise. These violations do not affect environment dynamics—they neither terminate episodes nor alter transitions. Importantly, our experiments do not access this cost directly during training; instead, the agent manages risk by being risk-aware in its velocities.
@@
Proximal Policy Optimization (PPO) [ppo] is used as a black-box solver to optimize the policy $\pi$, implementing line 4 of Algorithm 1. The dual variable $\lambda$ and auxiliary $t$ are updated via stochastic gradient steps following the rest of Algorithm 1. For further experimental details, see Appendix B.
@@
((a)) Environment: HalfCheetah   $c=1.450$   $\beta$-upper quantile: $1.419$   Converged $t$-value: $1.417$
((b)) Environment: Hopper   $c=0.373$   $\beta$-upper quantile: $0.370$   Converged $t$-value: $0.370$
((c)) Environment: Swimmer   $c=0.228$   $\beta$-upper quantile: $0.248$   Converged $t$-value: $0.207$
((d)) Environment: Walker2d   $c=1.171$   $\beta$-upper quantile: $1.133$   Converged $t$-value: $1.122$